match each equation to the corresponding graph
g 1 s=4sin (20) 2 = 6 cos e 3. r= 2 cos (30) 4. = 3 - 3 cos e 5. =4 -- 2 sin
Answer #1

match each Eguchion to the correspondivy graph -2 --- - -2 b) 2 1. r=4 sinao 2) 7=6 Coso 3) 7 = 20$(30) 4) r: 3-3 Coso 5) 7-4solutions -2 2- -- का Y(O)= |Qsino OSOM (3) 3- २(८(30) , O<< 10 -2 (0) -2 r=4 fin ao << 24 3-3-3500 6 ०%O.102 Y: 689 IC 2. OOT 지 나 C 2 =나- 28MO - 2 OCOST 니